### v2

@@ Abstract
Particle identification (PID) is essential for future particle physics experiments such as the Circular Electron-Positron Collider and the Future Circular Collider. A high-granularity Time Projection Chamber (TPC) not only provides precise tracking but also enables dNN/dxx measurements for PID. The dNN/dxx method estimates the number of primary ionization electrons, offering significant improvements in PID performance. However, accurate reconstruction remains a major challenge for this approach. In this paper, we introduce a deep learning model, the Graph Point Transformer (GraphPT), for dNN/dxx reconstruction. In our approach, TPC data are represented as point clouds. The network backbone adopts a U-Net architecture built upon graph neural networks, incorporating an attention mechanism for node aggregation specifically optimized for point cloud processing. The proposed GraphPT model surpasses the traditional truncated mean method in PID performance. In particular, the $K/\pi$ separation power improves by approximately 10% to 20% in the momentum interval from 5 to 20 GeV/$c$.
@@ 1 Introduction
Particle identification (PID) is a fundamental experimental technique in particle physics. At next-generation facilities such as the Circular Electron-Positron Collider (CEPC) cepc2018cepc ; gao2024cepc and the Future Circular Collider (FCC-ee) blondel2022fcc , hadron identification is expected to remain effective up to momenta of several tens of GeV/$c$. Conventional dEE/dxx measurements in current gaseous detectors besiii2009construction ; abe2010belle ; alme2010alice ; anderson2003star , which determine the total energy loss of a charged particle, cannot meet this requirement due to large fluctuations. Recent advances in electronics and detector technology now allow gaseous detectors to measure individual electron signals, enabling cluster counting (dNN/dxx). Unlike traditional dEE/dxx methods, the dNN/dxx method directly determines the number of primary ionization clusters davidenko1969tnmslation ; cataldi1997cluster , thereby suppressing the impact of secondary ionization as well as fluctuations from energy loss and amplification, and thus offering a significant enhancement in PID performance.
@@
The dNN/dxx method can be implemented in a Time Projection Chamber (TPC) when the readout granularity is sufficient to resolve individual electrons. Consequently, the baseline detector design for the CEPC features a large-volume, high-granularity TPC optimized for dNN/dxx measurements chang2023feasibility ; zhu2023requirement . Reconstruction, however, presents a significant challenge. In a large-volume TPC such as that proposed for the CEPC, the detector length is 5.8 m, corresponding to a maximum electron drift distance of 2.9 m. Such a long drift distance results in transverse diffusion on the order of several hundred micrometers, which complicates the use of spatial locality to distinguish overlapping clusters. Furthermore, the CEPC TPC has a pad size of $500\times 500\,\mu\textrm{m}^{2}$ to balance performance and power consumption. With this pad size and a gas mixture of Ar:CF4:iC4H10 (95:3:2), each pad collects on average about 1.8 primary electrons. Secondary electrons further increase this number, making dNN/dxx reconstruction particularly challenging.
@@
These challenges make traditional rule-based reconstruction algorithms highly difficult. Deep learning, a statistical-based approach that leverages large training datasets, can extract complex features from the data. In this paper, we propose a dNN/dxx reconstruction algorithm based on deep neural networks, which demonstrates superior performance compared with traditional methods. The remainder of this paper is organized as follows: Section 2 introduces the detector simulation framework and data samples; section 3 describes the traditional truncated mean reconstruction method; section 4 details the deep-learning–based reconstruction algorithm; section 5 presents the performances; and section 6 provides the conclusion and plan.
@@ 2.1 The CEPC TPC
Each endcap is equipped with 244 readout modules. Each module consists of a MicroMegas detector giomataris1996micromegas , a readout pad plane and the associated readout electronics. The modules also integrate all necessary high- and low-voltage supply cables, along with a cooling system to dissipate heat generated by the electronics.
A high-granularity readout liu2025development with a pad size of 500 $\times$ 500 $\rm\mu m^{2}$ was chosen based on simulation studies, with power consumption also taken into account. Each endcap contains more than 30 million readout channels in total. The fine granularity enables precise dNN/dxx measurements and has the potential to significantly enhance PID performance. The key design parameters of the CEPC TPC are summarized in table 1.
@@ 2.2 Simulation Framework
The TPC simulation framework is built on Garfield++ garfieldpp ; veenhof1993garfield with parameterizations that incorporate detector geometry, ionization processes, electron transport, electron amplification, and signal readout. The latter two stages are informed by results from existing prototype tests zhang2020study ; chang2025design ; liu2025development . As illustrated in figure 2, a full TPC chamber is implemented in the simulation. The gas mixture is set to Ar:CF4:iC4H10 (95:3:2). Ionization processes are simulated using the Heed code smirnov2005modeling , which is widely employed for precise ionization modeling in gases.
@@
The transported electrons are further amplified in the MicroMegas detector. The avalanche region of the MicroMegas provides a gain of 2000, and the single-electron gain distribution is modeled with a Polya function polya1930quelques . Within the framework, the number of electrons after amplification is sampled according to the Polya distribution, while the spatial density of the amplified electrons is described by a two-dimensional (2D) Gaussian distribution with an effective width $w$ (within $\pm 3\sigma$) of approximately 100 $\rm\mu m$.
@@ 2.3 Data Samples
Within the simulation framework, all pad-related hits are recorded for each simulated track. There is a one-to-one correspondence between hits and pads. For each hit, the deposited charge and timing information are stored, and its 2D position can be determined from the detector geometry. The charge and timing serve as input features for each hit, while the hit position is used to construct the point cloud representation. All input variables are normalized before being fed into the neural network.
@@ 3 Reconstruction with Truncated Mean Method
A conventional approach for suppressing secondary-electron effects is the truncated mean method, which has been widely applied in dEE/dxx reconstruction lippmann2012particle . In this approach, dEE/dxx—the energy deposition per unit track length—is used as the reconstruction variable. To mitigate resolution degradation caused by energy-loss fluctuations, typically $N$ independent dEE/dxx measurements $M$ are performed at multiple sampling points along the track within the TPC volume. For sufficiently large $N$, the resulting dEE/dxx distribution approaches a Landau distribution. Because of its long tail, the mean value of this distribution is not a reliable estimator of dEE/dxx. A more robust estimator is the truncated mean, in which the highest measurements—most likely associated with secondary electrons—are discarded. The truncated mean is defined as the average over the lowest $n$ values of the dEE/dxx measurements $M$:
@@
where $M_{i}\leq M_{i+1}$ for i=1,,N1i=1,...,N-1 and $\alpha=n/N$ is a fraction. The truncated mean dEE/dxx distribution is well approximated by a Gaussian. For dNN/dxx, the number of activated pads per unit track length serves as the ionization measurement, since it is approximately proportional to the number of ionization electrons. Consequently, the same truncated mean method can be applied.
@@ 4 Reconstruction with Deep Learning
Traditional rule-based approaches, such as the truncated mean method, rely heavily on human expertise and struggle to address the challenges outlined in section 1 for dNN/dxx reconstruction. In contrast, deep-learning–based methods can fully exploit detector information. A deep neural network, trained on large datasets, is capable of capturing hidden correlations among primary and secondary electrons as well as noise contributions. Therefore, we propose a deep-learning method, termed Graph Point Transformer (GraphPT), for dNN/dxx reconstruction.
4.1 Related Works
Self-attention and transformer.
Self-attention enables tokens to attend to all others, efficiently modeling long-range dependencies vaswani2017attention . Transformers stack self-attention and feed-forward layers for parallel sequence modeling, achieving state-of-the-art results in natural language processing and vision. Vision Transformers (ViT) extend this idea to images by treating patches as tokens, capturing both local and global structure, and outperforming Convolutional Neural Networks krizhevsky2012imagenet ; lecun1989backpropagation on large datasets dosovitskiy2020image ; kim2021rethinking .
Point-based networks.
Point-based networks adapt transformers to unordered three-dimensional (3D) point clouds by applying self-attention with positional encodings from coordinates, capturing both local geometry and global context. PointNet qi2017pointnet directly consumes point clouds while preserving permutation invariance. PointNet++ qi2017pointnet++ extends this with a hierarchical design that learns local features at multiple scales. Point Transformer (PT) zhao2021point replaces fixed pooling with adaptive self-attention, weighting neighbors by feature similarity and relative position to better capture local and long-range dependencies. Several improved PT variants have since been proposed wu2022point ; wu2024point .
Graph-based networks.
Graph Neural Networks (GNNs) scarselli2008graph model point clouds as graph nodes connected through neighborhood relations. This enables learning of both local geometry and global shape, making them effective for classification, segmentation, and object detection wang2019dynamic ; shi2020point ; li2024graph ; zheng2024pointvig .
dNN/dxx reconstruction.
@@ 4.1 Related Works
Traditional rule-based approaches to dNN/dxx reconstruction have been explored in several studies caron2014improved ; caputo2023particle ; aoki2022double . More recently, pioneering works have applied neural networks to dNN/dxx reconstruction, particularly in drift chambers, where signals are recorded as one-dimensional waveforms. Long Short-Term Memory networks and Dynamic Graph Convolutional Neural Networks (DGCNNs) wang2019dynamic have been applied to peak detection and clustering tasks with simulated data, achieving performance superior to traditional derivative-based and timing-based methods tian2025cluster . Furthermore, an innovative domain adaptation approach zhao2024peak , leveraging optimal transport theory kantorovich1942translocation , has been developed and successfully applied to test beam data.
@@ 4.2 Design Principle
For the TPC, each hit provides both 2D positional information and timing data, which together can be transformed into a 3D spatial coordinate. Consequently, a TPC track can be naturally represented as a point cloud. Inspired by the PT, we adopt a U-Net-based hierarchical architecture as the backbone. In this work, we further extend the framework by representing the irregularly distributed hits in 3D space as graphs, where the attention mechanism from the original PT is incorporated into the node aggregation process. This integration of GNNs allows more effective feature learning within the point cloud representation.
For the dNN/dxx reconstruction, each TPC hit is labeled either as positive or negative according to the number of primary electrons it contains, as determined from the MC truth information. Building on previous approaches, we streamline the reconstruction by unifying the two-step procedure described in ref. tian2025cluster into a single model. In this formulation, the reconstruction task can be defined as either a node classification problem or, equivalently, a point cloud segmentation problem.
@@ 4.3 Backbone Structure
As shown in figure 4, the backbone of the network is a U-Net-based ronneberger2015u hierarchical architecture, consisting of a U-shaped encoder–decoder with skip connections. To begin, a graph is constructed for each TPC track, where every TPC hit represented as a node. The node features include the hit’s charge and timing, and nodes are connected to their $k$ nearest neighbors (kkNN) based on Euclidean distance. The graph sequentially executes the hierarchical encoder and decoder.
Within the encoder, transitions down are performed by randomly dropping a fraction of nodes while projecting the features into a higher-dimension space. Conversely, in the decoder, transitions up are achieved through interpolation, with features mapped back into a lower-dimension space. At each transition, a new kkNN graph is constructed. Skip connections link layers at corresponding hierarchical levels, ensuring information flow across the network. Additionally, each transition layer incorporates a transformer layer, which will be discussed in the following section. At the end of the decoder, the output graph contains the same number of nodes as the initial input graph. Finally, an Multilayer Perceptron (MLP) layer followed by a softmax layer produces probabilities in the range [0, 1]. The detailed network parameters are summarized in table 2.
@@
Network Parameter GraphPT (Sub.) GraphPT (Dot-prod.) Feature Dimension Input 2 2 Encoder [32, 128, 512, 2048, 8192] [4, 16, 64, 256, 1024] Decoder [8192, 2048, 512, 128, 32] [1024, 256, 64, 16, 4] Output 1 1 Graph Construction $k$ in kkNN graph 16 16 Trans. down (up) ratio 0.25 (4) 0.25 (4) Transformer Layer Number of heads 8 Attention network $\delta$ MLP($d_{\textrm{out}}$, 64, $d_{\textrm{out}}$]) Positional network $\theta$ MLP([3, 64, $d_{\textrm{out}}$]) MLP([3, 64, 64, $d_{\textrm{out}}$])
@@ 4.4 Transformer Layer
The core component of the GraphPT architecture is the transformer layer. In the context of GNNs, the transformer layer performs message-passing mechanism: each node aggregates information from its neighbors to update its own representation. Formally, the message-passing can be expressed as:
@@
where $i$ denotes the index of the aggregation node with feature vector $\mathbf{x_{i}}$ and $\mathcal{N}(i)$ represents its neighboring nodes. The function $\beta$ maps the original feature vectors into a latent space, while the function $\alpha$ computes the attention weights used to aggregate neighbor information. The attention weights are defined as:
@@
The first operator is the subtraction operator that is used in PT zhao2021point , which is defined as:
@@
In contrast, the second operator is proposed in this work as a novel alternative. Specifically, the dot-product operator, which constitutes the fundamental mathematical component underlying self-attention, is adapted to the dNN/dxx reconstruction. It is defined as:

### v3

@@ Abstract
Particle identification (PID) is essential for future particle physics experiments such as the Circular Electron-Positron Collider (CEPC) and the Future Circular Collider. A high-granularity Time Projection Chamber (TPC) not only provides precise tracking but also enables dNN/dxx measurements for PID. The dNN/dxx method estimates the number of primary ionization electrons, offering significant improvements in PID performance. However, accurate reconstruction remains a major challenge for this approach. In this paper, we introduce a deep learning model, the Graph Point Transformer (GraphPT), for dNN/dxx reconstruction. In our approach, TPC data are represented as point clouds. The network backbone adopts a U-Net architecture built upon graph neural networks, incorporating an attention mechanism for node aggregation specifically optimized for point cloud processing. The proposed GraphPT model surpasses the traditional truncated mean method in PID performance. In particular, for the CEPC baseline TPC, the $K/\pi$ separation power improves by approximately 10% to 20% in the momentum interval from 5 to 20 GeV/$c$.
@@ 1 Introduction
Particle identification (PID) is a fundamental experimental technique in particle physics. At next-generation facilities such as the Circular Electron-Positron Collider (CEPC) Group and others (2018); Gao (2024) and the Future Circular Collider (FCC-ee) Blondel and Janot (2022), hadron identification is expected to remain effective up to momenta of several tens of GeV/$c$. Conventional dEE/dxx measurements in current gaseous detectors Collaboration and others (2009); Abe et al. (2010); Alme et al. (2010); Anderson et al. (2003), which determine the total energy loss of a charged particle, cannot meet this requirement due to large fluctuations. Recent advances in electronics and detector technology now enable gaseous detectors to achieve high resolution van Beuzekom et al. (2025, 2026); Group and others (2025); Amaro et al. (2023) and to measure individual electron signals, enabling cluster counting (dNN/dxx). Unlike traditional dEE/dxx methods, the dNN/dxx method directly determines the number of primary ionization clusters Davidenko et al. (1969); Cataldi et al. (1997), thereby suppressing the impact of secondary ionization as well as fluctuations from energy loss and amplification, and thus offering a significant enhancement in PID performance.
@@
The dNN/dxx method can be implemented in a Time Projection Chamber (TPC) when the readout granularity is sufficient to resolve individual electrons. Consequently, the baseline detector design for the CEPC features a large-volume, high-granularity TPC optimized for dNN/dxx measurements Qi and others (2024); Zhu et al. (2023). Reconstruction, however, presents a significant challenge. In a large-volume TPC such as that proposed for the CEPC, the detector length is 5.8 m, corresponding to a maximum electron drift distance of 2.9 m. Such a long drift distance results in transverse diffusion on the order of several hundred micrometers, which complicates the use of spatial locality to distinguish overlapping clusters. Furthermore, the CEPC TPC has a pad size of $500\times 500$ $\mu\textrm{m}^{2}$ to balance performance and power consumption. With this pad size and a gas mixture of Ar:CF4:iC4H10 (95:3:2), each pad that receives charge from a charged-particle track collects on average about 1.8 primary electrons. Secondary electrons further increase this number, making dNN/dxx reconstruction particularly challenging.
@@
These challenges make traditional rule-based reconstruction algorithms highly difficult. Deep learning, a statistics-based approach that leverages large training datasets, can extract complex features from the data. In this paper, we propose a dNN/dxx reconstruction algorithm based on deep neural networks, which demonstrates superior performance compared with traditional methods. The remainder of this paper is organized as follows: Section 2 introduces the detector simulation framework and data samples; section 3 describes the traditional truncated mean reconstruction method; section 4 details the deep-learning–based reconstruction algorithm; section 5 presents the performances; and section 6 provides the conclusion and plan.
@@ 2.1 The CEPC TPC
Each endcap is equipped with 244 readout modules. Each module consists of a MicroMegas detector Giomataris et al. (1996), a readout pad plane and the associated readout electronics. The modules also integrate all necessary high- and low-voltage supply cables, along with a cooling system to dissipate heat generated by the electronics.
A high-granularity readout Liu et al. (2025) with a pad size of 500 $\times$ 500 $\mu\textrm{m}^{2}$ was chosen based on simulation studies, with power consumption also taken into account. Each endcap contains more than 30 million readout channels in total. The fine granularity enables precise dNN/dxx measurements and has the potential to significantly enhance PID performance. The key design parameters of the CEPC TPC are summarized in table 1.
@@ 2.2 Simulation Framework
The TPC simulation framework is built on Garfield++ Schindler and Veenhof ; Veenhof and others (1993) with parameterizations that incorporate detector geometry, ionization processes, electron transport, electron amplification, and signal readout. The latter two stages are informed by results from existing prototype tests Zhang et al. (2020); Chang et al. (2025); Liu et al. (2025). As illustrated in figure 2, a full TPC chamber is implemented in the simulation. The gas mixture is set to Ar:CF4:iC4H10 (95:3:2). Ionization processes are simulated using the Heed code Smirnov (2005), which is widely employed for precise ionization modeling in gases.
@@
The transported electrons are further amplified in the MicroMegas detector. The avalanche region of the MicroMegas provides a gain of 2000, and the single-electron gain distribution is modeled with a Polya function Pólya (1930). Within the framework, the number of electrons after amplification is sampled according to the Polya distribution, while the spatial density of the amplified electrons is described by a two-dimensional (2D) Gaussian distribution. The effective width $w$ is defined as the the interval spanning $\pm 3\sigma$ of this 2D Gaussian—corresponding to the width illustrated in the right panel of figure 2—and is approximately 100 $\rm\mu m$.
@@ 2.3 Data Samples
Within the simulation framework, all pad-related hits are recorded for each simulated track. There is a one-to-one correspondence between hits and pads. For each hit, the deposited charge is stored, and its 2D ($x$, $y$) position is determined from the detector geometry. To obtain the $z$ position, the reconstructed track helix is first projected onto the $x$-$y$ plane, and the on the projected helix that lies closest to the TPC hit is identified as the “matched” point. The $z$ difference between the “matched” point and the hit is then taken as the hit’s $z$ coordinate, which can be converted to timing by dividing by the drift velocity. The charge and timing serve as input features for each hit, while the three-dimensional (3D) hit position ($x$, $y$, $z$) is used to construct the point cloud representation. All input variables are normalized before being fed into the neural network.
@@ 3 Reconstruction with Truncated Mean Method
A conventional approach for suppressing secondary-electron effects is the truncated mean method, which has been widely applied in dEE/dxx reconstruction Lippmann (2012). In this approach, dEE/dxx—the energy deposition per unit track length—is used as the reconstruction variable. To mitigate resolution degradation caused by energy-loss fluctuations, typically $N$ independent dEE/dxx measurements $M$ are performed at multiple sampling points along the track within the TPC volume. For sufficiently large $N$, the resulting dEE/dxx distribution approaches a Landau distribution. Because of its long tail, the mean value of this distribution is not a reliable estimator of dEE/dxx. A more robust estimator is the truncated mean, in which the highest measurements—most likely associated with secondary electrons—are discarded. The truncated mean is defined as the average over the lowest $n$ values of the dEE/dxx measurements $M$:
@@
where $M_{i}\leq M_{i+1}$ for i=1,,N1i=1,...,N-1 and $\alpha=n/N$ is a fraction. The truncated mean dEE/dxx distribution is well approximated by a Gaussian for sufficiently large $N$ and sufficiently small $\alpha$. For dNN/dxx, the number of activated pads per unit track length serves as the ionization measurement, since it is approximately proportional to the number of ionization electrons for sufficiently small pads. Consequently, the same truncated mean method can be applied.
@@ 4 Reconstruction with Deep Learning
Traditional rule-based approaches, such as the truncated mean method, rely on predefined selection rules and feature formulations, which limits their ability to address the challenges outlined in section 1 for dNN/dxx reconstruction. In contrast, deep-learning–based methods can fully exploit detector information. A deep neural network, trained on large datasets, is capable of capturing hidden correlations among primary and secondary electrons as well as noise contributions. Therefore, we propose a deep-learning method, termed Graph Point Transformer (GraphPT), for dNN/dxx reconstruction.
4.1 Related Work
The transition toward modern machine-learning in high-energy physics is driven by the need to resolve complex, non-linear patterns within the high-granularity data produced by next-generation detectors. A key breakthrough in this area is the self-attention mechanism Vaswani et al. (2017), which allows the algorithm to simultaneously evaluate the relationship between every individual signal “hit” and every other hit across the detector volume. This capability is crucial for capturing the global information of a particle event more effectively than local, rule-based algorithms Qu et al. (2022); Ratnikov and Rogachev (2021); Van Stroud et al. (2025).
To handle the spatial distribution of these hits, point-based architectures such as PointNet++ Qi et al. (2017a, b) introduce a hierarchical representation that enables the model to analyze fine-scale hit patterns while simultaneously learning broader track-level information. This approach can be further enhanced by the Point Transformer (PT) Zhao et al. (2021); Wu et al. (2022, 2024), which incorporates self-attention to dynamically weight neighboring hits based on their physical characteristics and relative spatial distances. In parallel, Graph Neural Networks (GNNs) Scarselli et al. (2008) treat detector data as a graph of interconnected nodes, where edges encode physical proximity. By learning over this graph structure, GNNs can effectively capture geometric shapes, neighborhood symmetries, and any other rational patterns intrinsic to particle interactions Duarte and Vlimant (2022); Qu and Gouskos (2020).
@@ 4.1 Related Work
These advancements are particularly relevant for dNN/dxx reconstruction. Beyond traditional rule-based approaches Caron et al. (2014); Caputo et al. (2023); Aoki et al. (2022), several pioneering studies have applied neural networks to dNN/dxx reconstruction, particularly in drift chambers, where signals are recorded as one-dimensional waveforms. Long Short-Term Memory networks and Dynamic Graph Convolutional Neural Networks (DGCNNs) Wang et al. (2019) have been applied to peak detection and clustering tasks with simulated data, achieving performance superior to traditional derivative-based and timing-based methods Tian et al. (2025). Furthermore, an innovative domain adaptation approach Zhao et al. (2024), leveraging optimal transport theory Kantorovich (1942), has been developed and successfully applied to test beam data.
@@ 4.2 Design Principle
For the TPC, each hit provides a 3D spatial coordinate. Consequently, a TPC track can be naturally represented as a point cloud. Inspired by the PT, a U-Net-based hierarchical architecture is adopted as the backbone. In this work, we further extend the framework by representing the irregularly distributed hits in 3D space as graphs, where the attention mechanism from the original PT is incorporated into the node aggregation process. This integration of GNNs allows more effective feature learning within the point cloud representation.
For the dNN/dxx reconstruction, each TPC hit is labeled either as positive or negative according to the number of primary electrons it contains, as determined from the MC truth information. Building on previous approaches, we streamline the reconstruction by unifying the two-step procedure described in ref. Tian et al. (2025) into a single model. In this formulation, the reconstruction task can be defined as either a node classification problem or, equivalently, a point cloud segmentation problem.
@@ 4.3 Backbone Structure
As shown in figure 4, the backbone of the network is a U-Net-based Ronneberger et al. (2015) hierarchical architecture, consisting of a U-shaped encoder–decoder with skip connections. To begin, a graph is constructed for each TPC track, where every TPC hit represented as a node. The node features include the hit’s charge and timing, and nodes are connected to their $k$ nearest neighbors (kkNN) based on 3D Euclidean distance. The graph sequentially executes the hierarchical encoder and decoder.
Within the encoder, transitions down are performed by applying farthest point sampling to a subset of nodes, while simultaneously projecting their features into a higher-dimension space. Conversely, in the decoder, transitions up are achieved through interpolation, with features mapped back into a lower-dimension space. At each transition, a new kkNN graph is constructed. Skip connections link layers at corresponding hierarchical levels, ensuring information flow across the network. Additionally, each transition layer incorporates a transformer layer, which will be discussed in the following section. At the end of the decoder, the output graph contains the same number of nodes as the initial input graph. Finally, an Multilayer Perceptron (MLP) layer followed by a softmax layer produces probabilities in the range [0, 1]. The detailed network parameters are summarized in table 2.
@@
Network Parameter GraphPT (Sub.) GraphPT (Dot-prod.) Feature Dimension Input 2 2 Encoder [32, 128, 512, 2048, 8192] [8, 32, 128, 512, 2048] Decoder [8192, 2048, 512, 128, 32] [2048, 512, 128, 32, 8] Output 1 1 Graph Construction $k$ in kkNN graph 16 16 Trans. down (up) ratio 0.25 (4) 0.25 (4) Transformer Layer Number of heads 4 Attention network $\delta$ MLP($d_{\textrm{out}}$, 64, $d_{\textrm{out}}$]) Positional network $\theta$ MLP([3, 64, $d_{\textrm{out}}$]) MLP([3, 64, 64, $d_{\textrm{out}}$])
@@ 4.4 Transformer Layer
The core component of the GraphPT architecture is the transformer layer. In the context of GNNs, the transformer layer performs message passing among nodes: each node receives information from its neighbors and aggregates it to update its own representation. The message-passing process can be formalized as:
@@
where $i$ denotes the index of the aggregation node with feature vector $\mathbf{x_{i}}$ and $\mathcal{N}(i)$ represents its neighboring nodes. The function $\beta$ is an MLP that maps the original feature vectors into a latent space, while the function $\alpha$ computes the attention weights used to aggregate neighbor information. The function $\alpha$ is defined as:
@@
The first operator is the subtraction operator that is used in PT Zhao et al. (2021), which is defined as:
@@
The dot-product operator, which is the fundamental mathematical component underlying self-attention, is further examined. It is defined as: